# Laplacian Spectrum and Domination in Trees

### v1

@@ 1 Introduction
We now list some of the known connections between $m_{G}I$ for different choices of $I$ and combinatorial parameters of graphs. In 1990, Grone, Merris, and Sunder grone1990laplacian proved that $m_{G}[0,1)\geqslant q(G)$, where $q(G)$ is the number of vertices in $G$ which are not pendant but adjacent to a pendant vertex. In 2013, Braga, Rodrigues, and Trevisan braga2013distribution demonstrated that for any tree $T$ with more than one vertex, the inequality $m_{T}[0,2)\geqslant\lceil\frac{n}{2}\rceil$ holds. They further conjectured that $m_{T}[0,\overline{d})\geqslant\lceil\frac{n}{2}\rceil$, where $\overline{d}$ represents the average degree of $T$. This conjecture was subsequently confirmed by Jacobs, Oliveira, and Trevisan jacobs2021most , and Sin sin2020number independently in 2021. In 2022, Ahanjideh et al. ahanjideh2022laplacian established that, for a connected graph $G$, $m_{G}[0,1)\leqslant\alpha(G)\leqslant m_{G}[0,n-\alpha(G)]$, where $\alpha(G)$ is the independence number of $G$. They also characterized graphs for which $m_{G}[0,1)=\alpha(G)=\frac{n}{2}$. Additionally, they discovered that for connected triangle-free or quadrangle-free graphs $G$, $m_{G}(n-1,n]\leqslant 1$, and characterized graphs such that $m_{G}[0,1]=n-1$ or n2n-2. Further, they showed that in connected graphs, $m_{G}(n-1,n]\leqslant\chi(G)-1$, where $\chi(G)$ is the chromatic number of $G$. Later, Akbari et al. akbari2023classification , and Cui and Ma cui2024laplacian derived several bounds on $m_{G}I$ based on various graph parameters. Also in 2024, Xu and Zhou explored the Laplacian eigenvalue distribution relative to the diameter of the graph xu2024diameter ; xu2024laplacian . Recently, Zhen, Wong, and Songnian Xu zhen2025laplacian and Leyou Xu and Zhou xu2025girth studied how the Laplacian eigenvalue distribution relates to the girth of a graph, while Akbari, Alaeiyan, and Darougheh akbari2025laplacian studied it in terms of degree sequence. In 2025, Guo, Xue and Liu guo2025laplacian proved that for a tree $T$ with diameter $d$, $m_{T}[0,1)\geqslant\frac{d+1}{3}$, and characterized trees that achieve this bound. Moon and Park moon2025laplacian gave a lower bound on $m_{G}[0,1)$ for unicyclic graphs in terms of its diameter and girth.
@@ 3.2 A New Dominating Set Algorithm
2:Initialize f(v):=d(v)1f(v):=d(v)-1 for all vertices $v$, $D=\emptyset$ and
$w(v)=\begin{cases}4/3,&\text{if $v$ is a penultimate vertex}\\ 0,&\text{otherwise}\end{cases}$
@@
10:   $w(v)=w(v)+1$
11:   if $v$ has a parent $u$, remove the edge vuvu
12:  end if
13:  if [($w(v)\geqslant 4/3$) or ($f(v)=0$ and $w(v)\geqslant 1$)] and at least one child of $v$ is not dominated by $D$ then
14:   $D=D\cup\{v\}$ and push a weight of (w(v)1w(v)-1) to $v$’s parent.
15:  else
16:   Push a weight of $w(v)$ to $v$’s parent
17:  end if
18:end for
19:return $D$
@@
We prove this by induction on the distance $d_{v}$ of $v$ to the nearest leaf below it. The base case is when $v$ is a penultimate vertex ($d_{v}=1)$. In this case, $v$ pushes a weight of 4/31=1/34/3-1=1/3 to its parent, confirming the claim. Let $v$ be a vertex with $d_{v}>1$ and let us assume that for every vertex $u$ with $d_{u}<d_{v}$, the claim is true. In particular, the claim is true for every child of $v$.
@@
For a vertex $v\in V(G)$, the notation $N[v]$ denotes the set of vertices adjacent to $v$, including $v$ itself. Since all the penultimate vertices are added to $D$, all the penultimates (which includes the root of $T$), and the leaves are dominated. Let $v$ be any non-root deep vertex of $T$ and $u$ be its parent. We will verify that the total weight entering $N[v]$ (including both initialization and propagation) is at least $4/3$, or at least $1$ with $f(u)=0$ when $u$ is being processed. This ensures the inclusion of at least one vertex of $N[v]$ in $D$, guaranteeing that $v$ is dominated by $D$.
@@ 4 Tight Example
Consider the tree $T_{k}$ illustrated in Figure 3. $T_{k}$, $k\geqslant 2$, has $k$ isomorphic subtrees hanging from the root. We only illustrate the leftmost subtree in the figure. By analyzing the run of Algorithm 1 on $T_{k}$, we can see that the negative vertices are the root vertex $r$ and the penultimate vertices. Hence $\mu(T_{k})=3k+1$. Furthermore, a minimum dominating set of $T_{k}$ must include at least four vertices from each subtree hanging from $r$. Hence $\gamma(T_{k})\geqslant 4k$. It is straightforward to construct a dominating set by picking four vertices from each subtree.
@@ 5 Improving the Upper Bound for High-Degree Trees
$w(v)=\begin{cases}1+\epsilon,&\text{if $v$ is a penultimate vertex}\\ 0,&\text{otherwise}\end{cases}$

### v2

@@ 1 Introduction
We now list some of the known connections between $m_{G}I$ for different choices of $I$ and combinatorial parameters of graphs. In 1990, Grone, Merris, and Sunder grone1990laplacian proved that $m_{G}[0,1)\geqslant q(G)$, where $q(G)$ is the number of vertices in $G$ which are not pendant but adjacent to a pendant vertex. In 2013, Braga, Rodrigues, and Trevisan braga2013distribution demonstrated that for any tree $T$ with more than one vertex, the inequality $m_{T}[0,2)\geqslant\lceil\frac{n}{2}\rceil$ holds. They further conjectured that $m_{T}[0,\overline{d})\geqslant\lceil\frac{n}{2}\rceil$, where $\overline{d}$ represents the average degree of $T$. This conjecture was subsequently confirmed by Jacobs, Oliveira, and Trevisan jacobs2021most, and Sin sin2020number independently in 2021. In 2022, Ahanjideh et al. ahanjideh2022laplacian established that, for a connected graph $G$, $m_{G}[0,1)\leqslant\alpha(G)\leqslant m_{G}[0,n-\alpha(G)]$, where $\alpha(G)$ is the independence number of $G$. They also characterized graphs for which $m_{G}[0,1)=\alpha(G)=\frac{n}{2}$. Additionally, they discovered that for connected triangle-free or quadrangle-free graphs $G$, $m_{G}(n-1,n]\leqslant 1$, and characterized graphs such that $m_{G}[0,1]=n-1$ or n2n-2. Further, they showed that in connected graphs, $m_{G}(n-1,n]\leqslant\chi(G)-1$, where $\chi(G)$ is the chromatic number of $G$. Later, Akbari et al. akbari2023classification, and Cui and Ma cui2024laplacian derived several bounds on $m_{G}I$ based on various graph parameters. Also in 2024, Xu and Zhou explored the Laplacian eigenvalue distribution relative to the diameter of the graph xu2024dia; xu2024laplacian. Recently, Zhen, Wong, and Songnian Xu zhen2025laplacian and Leyou Xu and Zhou xu2025girth studied how the Laplacian eigenvalue distribution relates to the girth of a graph, while Akbari, Alaeiyan, and Darougheh akbari2025laplacian studied it in terms of degree sequence. In 2025, Guo, Xue and Liu guo2025laplacian proved that for a tree $T$ with diameter $d$, $m_{T}[0,1)\geqslant\frac{d+1}{3}$, and characterized trees that achieve this bound. Moon and Park moon2025laplacian gave a lower bound on $m_{G}[0,1)$ for unicyclic graphs in terms of its diameter and girth.
@@ 3.2 A New Dominating Set Algorithm
2:Initialize f(v):=d(v)1f(v):=d(v)-1 for all vertices $v$, $D:=\emptyset$ and
$w(v):=\begin{cases}4/3,&\text{if $v$ is a penultimate vertex}\\ 0,&\text{otherwise}\end{cases}$
@@
10:   if $v$ has a parent $u$, remove the edge vuvu
11:  end if
12:  if $f(v)<0$ and $v$ is not a penultimate vertex then
13:   $w(v):=w(v)+1$
14:  end if
15:  if [($w(v)\geqslant 4/3$) or ($f(v)=0$ and $w(v)\geqslant 1$)] and at least one child of $v$ is not dominated by $D$ then
16:   $D:=D\cup\{v\}$ and push a weight of (w(v)1w(v)-1) to $v$’s parent.
17:  else
18:   Push a weight of $w(v)$ to $v$’s parent
19:  end if
20:end for
21:return $D$
@@
We prove this by induction on the distance $d_{v}$ of $v$ to the nearest leaf below it. The base case is when $v$ is a penultimate vertex ($d_{v}=1)$. In this case, $v$ pushes a weight of 4/31=1/34/3-1=1/3 to its parent, confirming the claim. Let $v$ be a vertex with $d_{v}>1$ and let us assume that for every vertex $x$ with $d_{x}<d_{v}$, the claim is true. In particular, the claim is true for every child of $v$.
@@
For a vertex $v\in V(G)$, the notation $N[v]$ denotes the set of vertices adjacent to $v$, including $v$ itself. Since all the penultimate vertices are added to $D$, all the penultimates (which includes the root of $T$), and the leaves are dominated. Let $v$ be any non-root deep vertex of $T$ and $u$ be its parent. We will verify that the total weight entering $N[v]$ (including both initialization and propagation) is at least $4/3$, or at least $1$ with $f(u)=0$ when $u$ is being processed. If no vertex in $N[v]\setminus\{u\}$ is added to $D$, then the total weight entering $N[v]$ reaches $u$. Since $u$ has an undominated child ($v$), and it satisfies the weight and $f(u)$ condition, $u$ will be added to $D$. This ensures the inclusion of at least one vertex of $N[v]$ in $D$, guaranteeing that $v$ is dominated by $D$.
@@ 4 Tight Example
Consider the tree $T_{k}$ illustrated in Figure 3. $T_{k}$, $k\geqslant 2$, has $k$ isomorphic subtrees hanging from the root. We only illustrate the leftmost subtree in the figure. By analyzing the run of Algorithm 1 on $T_{k}$, we can see that the negative vertices are the root vertex $r$ and the penultimate vertices. Hence $\mu(T_{k})=3k+1$. Furthermore, a minimum dominating set of $T_{k}$ must include at least four vertices from each subtree hanging from $r$. Hence $\gamma(T_{k})\geqslant 4k$. It is straightforward to construct a dominating set by picking four vertices from each subtree. Hence $\gamma(T_{k})=4k$ and
$\frac{\gamma(T_{k})}{\mu(T_{k})}=\frac{4k}{3k+1}.$
@@ 5 Improving the Upper Bound for High-Degree Trees
$w(v):=\begin{cases}1+\epsilon,&\text{if $v$ is a penultimate vertex}\\ 0,&\text{otherwise}\end{cases}$